# These $\altaffiliationThese authors contributed equally to this work.$

### v1

@@ Abstract
We introduce a family of all-electron Gaussian basis sets, augmented MOLOPT, optimized for excited-state calculations on large molecules and crystals. We generate these basis sets by augmenting existing STO-3G, STO-6G, and MOLOPT basis sets optimizied for ground state energy calculations. The augmented MOLOPT basis sets achieve fast convergence of GW gaps and Bethe–Salpeter excitation energies, while maintaining low condition numbers of the overlap matrix to ensure numerical stability. For GW HOMO-LUMO gaps, the double-zeta augmented MOLOPT basis yields a mean absolute deviation of 60 meV to the complete basis set limit. The basis set convergence for excitation energies from time-dependent density functional theory and the Bethe–Salpeter equation is similar. We use our smallest generated augmented MOLOPT basis (aug-SZV-MOLOPT-ae-mini) to demonstrate GW calculations on nanographenes with 2312 atoms requiring only 3500 core hours and 2.9 TB RAM as computational resources.
@@ 1 Introduction
First-principles electronic structure calculations 1 are now widely employed across various fields, including the computation of electronic band structures of crystals and the investigation of reaction mechanisms in chemistry. A fundamental initial step in nearly all such calculations is specifying the atomic geometry, that is, the positions of the atomic nuclei. While this task is relatively straightforward for small molecules or crystals with a few atoms in the unit cell, it quickly becomes complex as the number of atoms increases. For instance, determining the atomic geometry of a liquid-solid interface can be a challenge. Recently, there has been a transformative shift in how atomic geometries are determined, driven by advances in machine learning. Techniques such as machine-learned interatomic potentials 2 and direct structure prediction approaches, like those used in protein folding 3, are rapidly becoming standard tools in the field.
@@
As a result, increasingly complex atomic structures, comprising 100,000 atoms or more 4, are now available as starting points for first-principles calculations. A particularly interesting branch of these calculations is the study of electronically excited states 5, 6, 7. Understanding these excitations is important for interpreting optical experiments, from conventional optical absorption spectroscopy to ultrafast phenomena induced by femtosecond laser pulses 8. On the theoretical side, this poses a major challenge as first-principles methods for excited-state calculations are significantly more computationally demanding than those for ground-state properties 5, 6. The most widely used approaches for excited-state calculations include time-dependent density functional theory (TDDFT) 9, 10, the $GW$ approximation for quasiparticle energies, i.e., electron removal and addition energies, 11, 12, 13, and the $GW$ plus Bethe–Salpeter equation (GW-BSE) for optical excitations 5, 7. All of these methods have in common that the computational cost can quickly grow with the number of atoms in the molecule or unit cell, depending on the specific algorithm.
@@
One approach to restrict this growth in computation time is the usage of low-scaling algorithms which often employ spatial locality. As an example, we consider the irreducible density response $\chi^{0}(\mathbf{r},\mathbf{r}^{\prime})$, which describes how the electron density at position $\mathbf{r}$ changes in response to an external potential applied at position $\mathbf{r}^{\prime}$. $\chi^{0}(\mathbf{r},\mathbf{r}^{\prime})$ neglects the Coulomb interaction of the induced electron density and reflects the system’s intrinsic nonlocal polarizability. In semiconductors and dielectrics, $\chi^{0}(\mathbf{r},\mathbf{r}^{\prime})$ decays exponentially with increasing distance between $\mathbf{r}$ and $\mathbf{r}^{\prime}$, i.e., $\chi^{0}(\mathbf{r},\mathbf{r}^{\prime})\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt0$ as $|\mathbf{r}\hskip 1.69998pt{-}\hskip 1.69998pt\mathbf{r}^{\prime}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$, which is a sign of the spatial locality of the electronic structure, the "nearsightedness" 14, 15.
@@
Spatial locality can be exploited in the $GW$ space-time algorithm 16, 17, which scales as $O(N^{3})$ with the number of atoms $N$, compared to the $O(N^{4})$ scaling of conventional frequency-based $GW$ algorithms 13. The space-time algorithm uses a real-space grid representation and switches between real-space and plane-wave bases. To retain the favorable scaling, key steps must be performed in real space; using plane waves throughout would increase scaling again to $O(N^{4})$. Fast Fourier transforms (FFTs) enable efficient switching between representations but require an equidistant real-space grid. Such grids are not ideal for electronic structure calculations, as they fail to exploit the characteristic shapes of atomic orbitals, such as $s$-, $p$-, or $d$-type functions 111 We mention recent advancements in the generation of non-equidistant real-space grids 18, 19, 20, 21, 22 which have reduced the number of grid points dramatically. Still, a compact atomic-orbital basis set is required for the generation of these real-space grids.. Reformulating the $GW$ space-time method in a localized atomic-orbital basis can significantly reduce matrix sizes, enabling efficient calculations for twodimensional crystals 23 as well as large and complex systems 24, 25, 21, 26, 27.
@@
The computational effort of excited-state methods like $GW$, TDDFT and GW-BSE depends sensitively on the size of the atomic-orbital basis set. As an example, the computational cost of space-time $GW$ 24, 26, 27 increases with the fourth power in the number of basis functions per atom. It is thus highly desirable to employ optimal atomic-orbital basis sets that provide converged excitation energies with a small number of basis functions. Atomic-orbital basis sets come in various forms, including numeric atom-centered orbitals (NAOs)28, 29, Slater-type orbitals (STOs)30, and Gaussian-type orbitals (GTOs)31. NAOs offer high flexibility and accuracy with compact sets, while STOs closely resemble atomic orbitals and also allow for small basis sizes. GTOs achieve radial flexibility via superposition of several Gaussians ("contractions") and enable efficient evaluation of Coulomb integrals through analytical expressions 31, 32, 33. This efficiency has made GTOs a standard in quantum chemistry software and motivates their use in this work.
@@
Most GTO basis sets have been optimized for the computation of the ground state energy. However, when such basis sets are used in excited-state methods like GW-BSE, they yield a slow basis-set convergence. To cure this issue, one can add further Gaussian functions to ground-state optimized basis sets, as it is done in the aug-cc-pVXZ basis set family, X = D, T, Q, 5 34. Here, the additional Gaussians are optimized to match the LUMO wavefunction via optimization of the total energy of charged atoms. In this way, the additional Gaussians describe electronically excited states, which often involve the excitation from occupied orbitals to LUMO. The aug-cc-pVXZ basis sets yield good accuracy for excitation energies of small molecules, but their application to large or periodic systems is severely limited by numerical issues mainly related to the inclusion of very diffuse Gaussian functions, i.e., those with very small exponents that decay slowly. As a result, the condition number of the overlap matrix is large 35, leading to numerical instability and convergence problems in the self-consistent-field iterations of solids and large molecules. Consequently, the use of aug-cc-pVXZ is typically prohibitive for large molecules and solids. This presents a gap in the current methodology: while Gaussian basis sets optimized for excited-state calculations of small molecules exist, Gaussian basis sets optimized for excited-state calculations of solids and large, complex molecular systems are lacking.
@@
In contrast to excited states, Gaussian basis sets tailored for computing the DFT ground-state energy in solids and large molecular systems already exist. A prominent class of such basis sets are the MOLOPT-type basis sets, which were specifically designed to balance accuracy of DFT ground-state energy calculations with numerical stability. Not only the accuracy of ground-state energies of molecules 35, 36, 37, 38 has been optimized, but also the condition number of the overlap matrix has been minimized, which is critical for ensuring numerically stable DFT calculations in extended and complex systems.
@@
Despite their success in ground-state calculations, the currently available MOLOPT basis sets 35, 36, 37, 38 do not offer a sufficiently accurate description of electronically excited states. In this work, we address this limitation by augmenting existing all-electron MOLOPT basis sets 37, 38 with additional diffuse Gaussian functions, which are optimized to reproduce excitation energies from $GW$-BSE calculations performed with a large reference basis set (aug-cc-pV5Z). As a result, we introduce the all-electron (ae) aug-MOLOPT-ae basis set familiy containing aug-SZV-MOLOPT-ae, aug-DZVP-MOLOPT-ae, and aug-TZVP-MOLOPT-ae. Our basis sets cover the elements of periods I, II, and III (H to Cl) and are specifically designed for accurate and efficient excited-state calculations in large molecular systems and the condensed phase.
@@
The article is organized as follows: In Sec. 2, we give an overview of Kohn-Sham DFT in a Gaussian basis, where numerical instabilities due to the inverse overlap matrix can arise, as discussed in detail in Sec. 3. A theoretical perspective on basis set convergence for quasiparticle and excitation energies is provided in Sec. 4. We further provide the procedure for generating the augmented MOLOPT basis sets in Sec. 5 and associated auxiliary RI basis sets in Sec. 6. We describe the molecular test set for our benchmark in Sec. 7 and the computational details in Sec. 8. Benchmark results on HOMO-LUMO gaps from PBE0 and $GW$ as well as excitation energies from BSE and TDDFT for the augmented MOLOPT basis sets are given in Sec. 9 and 10, respectively. In Sec. 11, we evaluate how auxiliary RI basis size and Coulomb cutoff affect the accuracy in low-scaling $GW$ calculations. Large-scale applicability of the augmented MOLOPT basis sets is demonstrated in Sec. 12, where we perform $GW$ calculations on nanographenes with over 2000 atoms.
2 Expansion of Kohn-Sham orbitals
@@
Many excited-state calculations of solids and large molecules start from Kohn-Sham (KS) density functional theory (or Hartree-Fock theory) 39, 40, where one needs to solve the KS equations,
@@
To solve the KS equations (1) numerically, each KS orbital is expanded as a linear combination of predefined basis functions. When a Gaussian basis set is used, each orbital $\psi_{n}(\mathbf{r})$ is written as a sum over basis functions $\phi_{\mu}(\mathbf{r})$ of the molecule (or unit cell 29) with expansion coefficients $C_{\mu n}$,
@@
known as the Roothaan-Hall equations 40. Here, $h_{\mu\nu}$ is the Kohn-Sham matrix and the overlap matrix is defined as
@@
Eq. (3) is a generalized eigenvalue problem because the basis functions $\{\phi_{\mu}\}$ are non-orthogonal for molecules with more than a single atom, i.e., $\mathbf{S}\hskip 1.69998pt{\neq}\hskip 1.69998pt\mathbf{Id}$. To solve Eq. (3) for $C_{\nu n}$ and  $\varepsilon_{n}$, one usually transforms Eq. (3) into a standard eigenvalue problem
@@
A Gaussian-type basis function $\phi_{\mu}(\mathbf{r})$ used in Eq. (2) is centered at an atom $A$ and is a linear combination of Gaussian functions multiplied with a spherical harmonic $Y_{l}^{m}$41, 40
@@
3 Numerical instability computing $\mathbf{S}^{\text{--1/2}}$ and the condition number of $\mathbf{S}$
While Gaussian basis sets offer powerful flexibility by tuning contraction coefficients $\alpha$ and exponents $\beta$, Eq. (7), the inclusion of diffuse Gaussian functions, i.e., those with small exponents, can introduce serious numerical challenges. These diffuse functions are often essential for accurately capturing excited-state quantities, because empty KS orbitals are usually more diffuse than occupied KS orbitals. Diffuse Gaussians decay slowly and exhibit significant spatial overlap even across distant atoms. As a result, the overlap matrix $\mathbf{S}$ becomes increasingly ill-conditioned, with eigenvalues that span several orders of magnitude. This poor conditioning leads to numerical instability when computing $\mathbf{S}^{-1/2}$ for transforming the generalized KS eigenvalue problem (3) into a standard one (6). Such instabilities are particularly challenging in large molecules or periodic systems which can lead to convergence issues of the self-consistent-field (SCF) cycle. In the following, we analyze this numerical instability in detail.
@@ 3 Numerical instability computing $\mathbf{S}^{\text{--1/2}}$ and the condition number of $\mathbf{S}$
Numerical instabilities arise if $\kappa(\mathbf{S})$ hits the inverse machine precision. For double precision arithmetic, machine precision is $2^{-52}\hskip 1.69998pt{\approx}\hskip 1.69998pt10^{-16}$ and thus numerical instabilities arise if $\kappa(\mathbf{S})\gtrsim 10^{16}$. Several numerical tricks have been used to circumvent these instabilities related to large $\kappa(\mathbf{S})$, for example the removal of small eigenvalues of the overlap matrix 35. In our experience, we have observed that this technique can help to certain extent, but for too large condition numbers, the SCF cycle fails to converge nevertheless.
@@
Figure 1: Condition number $\kappa(\mathbf{S})$ of the overlap matrix for a finite graphite-like cutout computed from Eq. (11) using five different Gaussian basis sets. The cutout with 1430 atoms is shown in the inset; we vary its horizontal length, and the corresponding number of atoms is plotted on the horizontal axis. For the aug-cc-pVDZ basis set 42, 43, $\kappa(\mathbf{S})$ exceeds the inverse machine precision ($\sim 10^{16}$) for cutouts with 750 atoms and more, leading convergence issues in the SCF. We also show the condition number of the four augmented MOLOPT basis sets (aug-SZV-MOLOPT-ae-mini, aug-SZV-MOLOPT-ae-SR, aug-SZV-MOLOPT-ae, aug-DZVP-MOLOPT-ae) developed in this work, where the condition number remains well below this threshold.
$\kappa(\mathbf{S})$ increases rapidly with additional diffuse functions per atom and for systems with more atoms. For example, when considering a crystal, the condition number of the overlap matrix $\mathbf{S}(\mathbf{k})$ 28, 29
$S_{\mu\nu}(\mathbf{k})=\sum_{\mathbf{R}}e^{i\mathbf{k}\cdot\mathbf{R}}\int_{\mathbb{R}^{3}}d\mathbf{r}\;\phi_{\mu}(\mathbf{r})\,\phi_{\nu}(\mathbf{r}-\mathbf{R})$ (12)
@@
can be huge if diffuse Gaussian are contained in the basis set. Here, $\mathbf{k}$ is the crystal momentum from the first Brillouin zone and $\mathbf{R}$ are lattice vectors. We illustrate the effect of the lattice sum (12) on the condition number (11) in Fig. 1: When adding more atoms to the system, $\kappa(\mathbf{S})$ increases by orders of magnitude for the commonly used Gaussian aug-cc-pVDZ basis set (dark green curve) 42, 43, reaching values above the inverse machine precision, $\kappa(\mathbf{S})\hskip 1.69998pt{>}\hskip 1.69998pt10^{16}$. In contrast, for all four augmented MOLOPT basis sets presented later in this work, the condition number remains well below this threshold, even in the limit of infinite system size (bulk graphite).
4 Basis set convergence for excitation energies of charged and neutral excitations

### v2

@@ Abstract
We introduce a family of all-electron Gaussian basis sets, augmented MOLOPT, optimized for excited-state calculations on large molecules. We generate these basis sets by augmenting existing STO-3G, STO-6G, and MOLOPT basis sets optimized for ground state energy calculations. The augmented MOLOPT basis sets achieve fast convergence of GW gaps and Bethe–Salpeter excitation energies, while maintaining low condition numbers of the overlap matrix to ensure numerical stability. For GW HOMO-LUMO gaps, the double-zeta augmented MOLOPT basis yields a mean absolute deviation of 60 meV to the complete basis set limit. The basis set convergence for excitation energies from time-dependent density functional theory and the Bethe–Salpeter equation is similar. We use our smallest generated augmented MOLOPT basis (aug-SZV-MOLOPT-ae-mini) to demonstrate GW calculations on nanographenes with 9224 atoms requiring only 34300 core hours of computational resources.
@@ 1 Introduction
First-principles electronic structure calculations Martin2004 are now widely employed across various fields, including the computation of electronic band structures of crystals and the investigation of reaction mechanisms in chemistry. A fundamental initial step in nearly all such calculations is specifying the atomic geometry, that is, the positions of the atomic nuclei. While this task is relatively straightforward for small molecules with a few atoms in the unit cell, it quickly becomes complex as the number of atoms increases. For instance, determining the atomic geometry of a liquid-solid interface can be a challenge. Recently, there has been a transformative shift in how atomic geometries are determined, driven by advances in machine learning. Techniques such as machine-learned interatomic potentials Behler2016 and direct structure prediction approaches, like those used in protein folding Jumper2021, are rapidly becoming standard tools in the field.
@@
As a result, increasingly complex atomic structures, comprising 100,000 atoms or more Deringer2021, are now available as starting points for first-principles calculations. A particularly interesting branch of these calculations is the study of electronically excited states Onida2002, Dreuw2005, Blase2018. Understanding these excitations is important for interpreting optical experiments, from conventional optical absorption spectroscopy to ultrafast phenomena induced by femtosecond laser pulses Krasuz2009. On the theoretical side, this poses a major challenge as first-principles methods for excited-state calculations are significantly more computationally demanding than those for ground-state properties Onida2002, Dreuw2005. The most widely used approaches for excited-state calculations include time-dependent density functional theory (TDDFT) Runge1984, Ullrich2011, the $GW$ approximation for quasiparticle energies, i.e., electron removal and addition energies hedin1965new, Reining2018, Golze2019, and the $GW$ plus Bethe–Salpeter equation (GW-BSE) for optical excitations Onida2002, Blase2018. All of these methods have in common that the computational cost can quickly grow with the number of atoms in the molecule or unit cell, depending on the specific algorithm.
@@
One approach to restrict this growth in computation time is the usage of low-scaling algorithms which often employ spatial locality. As an example, we consider the irreducible density response $\chi^{0}(\mathbf{r},\mathbf{r}^{\prime})$, which describes how the electron density at position $\mathbf{r}$ changes in response to an external potential applied at position $\mathbf{r}^{\prime}$. $\chi^{0}(\mathbf{r},\mathbf{r}^{\prime})$ neglects the Coulomb interaction of the induced electron density and reflects the system’s intrinsic nonlocal polarizability. In semiconductors and dielectrics, $\chi^{0}(\mathbf{r},\mathbf{r}^{\prime})$ decays exponentially with increasing distance between $\mathbf{r}$ and $\mathbf{r}^{\prime}$, i.e., $\chi^{0}(\mathbf{r},\mathbf{r}^{\prime})\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt0$ as $|\mathbf{r}\hskip 1.69998pt{-}\hskip 1.69998pt\mathbf{r}^{\prime}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$, which is a sign of the spatial locality of the electronic structure, the "nearsightedness" Kohn1996, Prodan2005.
@@
Spatial locality can be exploited in the $GW$ space-time algorithm rojas_1995, liu2016, which scales as $O(N^{3})$ with the number of atoms $N$, compared to the $O(N^{4})$ scaling of conventional frequency-based $GW$ algorithms Golze2019. The space-time algorithm uses a real-space grid representation and switches between real-space and plane-wave bases. To retain the favorable scaling, key steps must be performed in real space; using plane waves throughout would increase scaling again to $O(N^{4})$. Fast Fourier transforms (FFTs) enable efficient switching between representations but require an equidistant real-space grid. Such grids are not ideal for electronic structure calculations, as they fail to exploit the characteristic shapes of atomic orbitals, such as $s$-, $p$-, or $d$-type functions. We do need to mention recent advancements in the generation of nonequidistant real-space grids Lu2015, Lu2017, Duchemin2019, duchemin2021, Delesma2024 which have reduced the number of grid points dramatically. Still, a compact atomic-orbital basis set is required for the generation of these real-space grids. Reformulating the $GW$ space-time method in a localized atomic-orbital basis can significantly reduce matrix sizes, enabling efficient calculations for two-dimensional crystals Pasquier2025 as well as large and complex systems wilhelm2018, foerster2020, duchemin2021, foerster2022quasiparticle, Graml2024.
@@
The computational effort of excited-state methods like $GW$, TDDFT and GW-BSE depends sensitively on the size of the atomic-orbital basis set. As an example, the computational cost of space-time $GW$ wilhelm2018, foerster2022quasiparticle, Graml2024 increases with the fourth power in the number of basis functions per atom. It is thus highly desirable to employ optimal atomic-orbital basis sets that provide converged excitation energies with a small number of basis functions. Atomic-orbital basis sets come in various forms, including numeric atom-centered orbitals (NAOs)Delley2000, Blum2009, Slater-type orbitals (STOs)tevelde2001, and Gaussian-type orbitals (GTOs)hehre1969a. NAOs offer high flexibility and accuracy with compact sets, while STOs closely resemble atomic orbitals and also allow for small basis sizes. GTOs achieve radial flexibility via superposition of several Gaussians ("contractions") and enable efficient evaluation of Coulomb integrals through analytical expressions hehre1969a, Obara1986, Golze2017. This efficiency has made GTOs a standard in quantum chemistry software and motivates their use in this work.
@@
Most GTO basis sets have been optimized for the computation of the ground state energy. However, when such basis sets are used in excited-state methods like GW-BSE, they yield a slow basis-set convergence. To cure this issue, one can add further Gaussian functions to ground-state optimized basis sets, as it is done in the aug-cc-pVXZ basis set family, X = D, T, Q, 5 kendall1992a. Here, the additional Gaussians are optimized to match the LUMO wave function via optimization of the total energy of charged atoms. In this way, the additional Gaussians describe electronically excited states, which often involve the excitation from occupied orbitals to LUMO. The aug-cc-pVXZ basis sets yield good accuracy for excitation energies of small molecules, but their application to large systems is severely limited by numerical issues mainly related to the inclusion of very diffuse Gaussian functions, i.e., those with very small exponents that decay slowly. As a result, the condition number of the overlap matrix is large Vandevondele2007, leading to numerical instability and convergence problems in the self-consistent-field iterations of large molecules. Consequently, the use of aug-cc-pVXZ is typically prohibitive for large molecules . This presents a gap in the current methodology: while Gaussian basis sets optimized for excited-state calculations of small molecules exist, Gaussian basis sets optimized for excited-state calculations of large, complex molecular systems are lacking.
@@
In contrast to excited states, Gaussian basis sets tailored for computing the DFT ground-state energy in large molecular systems already exist. A prominent class of such basis sets are the MOLOPT-type basis sets, which were specifically designed to balance accuracy of DFT ground-state energy calculations with numerical stability. Not only the accuracy of ground-state energies of molecules Vandevondele2007, YeBerkelbach2022, MOLOPT_PBE_ae, PhDTMueller has been optimized, but also the condition number of the overlap matrix has been minimized, which is critical for ensuring numerically stable DFT calculations in extended and complex systems.
@@
Despite their success in ground-state calculations, the currently available MOLOPT basis sets Vandevondele2007, YeBerkelbach2022, MOLOPT_PBE_ae, PhDTMueller do not offer a sufficiently accurate description of electronically excited states. In this work, we address this limitation by augmenting existing all-electron MOLOPT basis sets MOLOPT_PBE_ae, PhDTMueller with additional diffuse Gaussian functions, which are optimized to reproduce excitation energies from $GW$-BSE calculations performed with a large reference basis set (aug-cc-pV5Z). As a result, we introduce the all-electron (ae) aug-MOLOPT-ae basis set family containing aug-SZV-MOLOPT-ae, aug-DZVP-MOLOPT-ae, and aug-TZVP-MOLOPT-ae. Our basis sets cover the elements of periods I, II, and III (H to Cl) and are specifically designed for accurate and efficient excited-state calculations in large molecular systems .
@@
The article is organized as follows: In Sec. 2.1, we give an overview of Kohn-Sham DFT in a Gaussian basis, where numerical instabilities due to the inverse overlap matrix can arise, as discussed in detail in Sec. 2.2. A theoretical perspective on basis set convergence for quasiparticle and excitation energies is provided in Sec. 2.3. We further provide the procedure for generating the augmented MOLOPT basis sets in Sec. 2.4 . Benchmark results on HOMO-LUMO gaps from PBE0 and $GW$ as well as excitation energies from BSE and TDDFT for the augmented MOLOPT basis sets are given in Sec. 2.5 and 2.6, respectively. The procedure to generate the associated auxiliary RI basis sets is presented in Sec. 3.1. In Sec. 3.2, we evaluate how auxiliary RI basis size and Coulomb cutoff affect the accuracy in low-scaling $GW$ calculations. Large-scale applicability of the augmented MOLOPT basis sets is demonstrated in Sec. 4, where we perform $GW$ calculations on nanographenes with over 9000 atoms. We describe the molecular test set for our benchmark and the computational details in App. A. In App. B, we describe a memory-saving scheme for the computation of the self-energy using a repeated calculation of three-centre integrals. Finally, in App. C, we carry out a test calculation of our new basis sets on a representative system (9,10-Dihydroanthracene), in order to quantitatively assess the quality of these basis sets with respect to their size.
2 Orbital basis sets
2.1 Expansion of Kohn-Sham orbitals
@@
Many excited-state calculations of large molecules start from Kohn-Sham (KS) density functional theory (or Hartree-Fock theory) Kohn1965, jensen, where one needs to solve the KS equations,
@@
To solve the KS equations (1) numerically, each KS orbital is expanded as a linear combination of predefined basis functions. When a Gaussian basis set is used, each orbital $\psi_{n}(\mathbf{r})$ is written as a sum over basis functions $\phi_{\mu}(\mathbf{r})$ of the molecule (or unit cell Blum2009) with expansion coefficients $C_{\mu n}$,
@@
known as the Roothaan-Hall equations jensen. Here, $h_{\mu\nu}$ is the Kohn-Sham matrix and the overlap matrix is defined as
@@
Eq. (3) is a generalized eigenvalue problem because the basis functions $\{\phi_{\mu}\}$ are nonorthogonal for molecules with more than a single atom, i.e., $\mathbf{S}\hskip 1.69998pt{\neq}\hskip 1.69998pt\mathbf{Id}$. To solve Eq. (3) for $C_{\nu n}$ and  $\varepsilon_{n}$, one usually transforms Eq. (3) into a standard eigenvalue problem
@@
A Gaussian-type basis function $\phi_{\mu}(\mathbf{r})$ used in Eq. (2) is centered at an atom $A$ and is a linear combination of Gaussian functions multiplied with a spherical harmonic $Y_{l}^{m}$Kuehne2020, jensen
@@
2.2 Numerical instability computing $\mathbf{S}^{\text{--1/2}}$ and the condition number of $\mathbf{S}$
While Gaussian basis sets offer powerful flexibility by tuning contraction coefficients $\alpha$ and exponents $\beta$, Eq. (7), the inclusion of diffuse Gaussian functions, i.e., those with small exponents, can introduce serious numerical challenges. These diffuse functions are often essential for accurately capturing excited-state quantities, because empty KS orbitals are usually more diffuse than occupied KS orbitals. Diffuse Gaussians decay slowly and exhibit significant spatial overlap even across distant atoms. As a result, the overlap matrix $\mathbf{S}$ becomes increasingly ill-conditioned, with eigenvalues that span several orders of magnitude. This poor conditioning leads to numerical instability when computing $\mathbf{S}^{-1/2}$ for transforming the generalized KS eigenvalue problem (3) into a standard one (6). Such instabilities are particularly challenging in large molecules which can lead to convergence issues of the self-consistent-field (SCF) cycle. In the following, we analyze this numerical instability in detail.
@@ 2.2 Numerical instability computing $\mathbf{S}^{\text{--1/2}}$ and the condition number of $\mathbf{S}$
Numerical instabilities arise if $\kappa(\mathbf{S})$ hits the inverse machine precision. For double precision arithmetic, machine precision is $2^{-52}\hskip 1.69998pt{\approx}\hskip 1.69998pt10^{-16}$ and thus numerical instabilities arise if $\kappa(\mathbf{S})\gtrsim 10^{16}$. Several numerical tricks have been used to circumvent these instabilities related to large $\kappa(\mathbf{S})$, for example the removal of small eigenvalues of the overlap matrix Vandevondele2007. In our experience, we have observed that this technique can help to certain extent, but for too large condition numbers, the SCF cycle fails to converge nevertheless. One of the possible reasons for this behaviour is that the eigendecomposition of the overlap matrix can become highly unstable for overcomplete basis sets Lehtola2019, leading to an unreliable regularization and thus SCF calculation.
@@
Figure 1: Condition number $\kappa(\mathbf{S})$ of the overlap matrix for a finite graphite-like cutout computed from Eq. (11) using five different Gaussian basis sets. The cutout with 1430 atoms is shown in the inset; we vary its horizontal length, and the corresponding number of atoms is plotted on the horizontal axis. For the aug-cc-pVDZ basis set dunning1989a, woon1994a, $\kappa(\mathbf{S})$ exceeds the inverse machine precision ($\sim 10^{16}$) for cutouts with 750 atoms and more, leading to convergence issues in the SCF. We also show the condition number of the four augmented MOLOPT basis sets (aug-SZV-MOLOPT-ae-mini, aug-SZV-MOLOPT-ae-SR, aug-SZV-MOLOPT-ae, aug-DZVP-MOLOPT-ae) developed in this work, where the condition number remains well below this threshold.
@@
$\kappa(\mathbf{S})$ increases rapidly with additional diffuse functions per atom and for systems with more atoms. We illustrate the effect of the molecular size on the condition number (11) in Fig. 1: When adding more atoms to the system, $\kappa(\mathbf{S})$ increases by orders of magnitude for the commonly used Gaussian aug-cc-pVDZ basis set (dark green curve) dunning1989a, woon1994a, reaching values above the inverse machine precision, $\kappa(\mathbf{S})\hskip 1.69998pt{>}\hskip 1.69998pt10^{16}$. In contrast, for all four augmented MOLOPT basis sets presented later in this work, the condition number remains well below this threshold, even in the limit of infinite system size (bulk graphite).
2.3 Basis set convergence for excitation energies of charged and neutral excitations